OpenStudy (anonymous):

3(x + 1) = -2(x - 1) + 6

OpenStudy (rulnick):

3x + 3 = -2x + 2 + 6 3x + 3 = -2x + 8 5x = 5 x=1

OpenStudy (anonymous):

3(x+1)=-2(x-1)+6= 3x+3=-2x+2+6= 3x+3=-2x+8= -3 -3= 3x=-2x+8= +2x +2x 5x=5 _______ 5 5 x=1